AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 341     Entry time: Tue Mar 1 05:23:02 2022
Author: JM, EM 
Subject: Current status  

06:20

Servo motors moved to fully out position as some work on the beam was neccessary. Beam injection had dropped from ~9e7 to around 1e7. Potentiometer reads 19.10 mm.   

Still trying to improve target density. Current setpoint 100 K.

Merger OK

Data rate = 1600 kb/sec

Leakage current = 3.21 uA

System checks OK - wr fails

         Base         Current     Difference
aida01 fault      0xab60 :      0xab8e :      46  
aida02 fault      0x88d3 :      0x8900 :      45  
aida03 fault      0x23aa :      0x23d8 :      46  
aida04 fault      0x15d7 :      0x1605 :      46  
aida05 fault      0x2f03 :      0x2f0f :      12  
White Rabbit error counter test result: Passed 0, Failed 5

Understand the status reports as follows:-
Status bit 3 : White Rabbit decoder detected an error in the received data
Status bit 2 : Firmware registered WR error, no reload of Timestamp
Status bit 0 : White Rabbit decoder reports uncertain of Timestamp information from WR

Avg ring pressure = 4.3E-10 mbar.

06:40

Target density still very low. Setpoint back to 116 K. Detectors still fully out, very low rutherford rate.

Attempted to move nozzle to increase density, already on maximum point. 

06:50

Setpoint to 95 k. Small increase in target density, however this was temporary

 

Attachment 1: 2022-03-01_06-42-53-ExRate.png  157 kB  | Hide | Hide all
2022-03-01_06-42-53-ExRate.png
ELOG V3.1.4-unknown